### v1

@@ Abstract
Methods: We developed an eight-agent architecture with centralized control mechanisms for dynamic medical consultation. The framework decomposes pre-consultation into four primary tasks: Triage ($T_{1}$), History of Present Illness collection ($T_{2}$), Past History collection ($T_{3}$), and Chief Complaint generation ($T_{4}$), with $T_{1}$$T_{3}$ further divided into 13 domain-specific subtasks. A Controller agent coordinates specialized agents through dynamic completion assessment, adaptive prompt generation, and hierarchical task management. The system was evaluated on 1,372 validated electronic health records from a Chinese medical platform, with performance assessed across multiple foundation models (GPT-OSS 20B, Qwen3-8B, Phi4-14B).
Results: The framework achieved 87.0% accuracy for primary department triage and 80.5% for secondary department classification. Task completion rates reached 98.2% using agent-driven scheduling versus 93.1% with sequential processing. Clinical quality scores from 18 physicians averaged 4.56 for Chief Complaints, 4.48 for History of Present Illness, and 4.69 for Past History (5-point scale). The system demonstrated robust performance across different language models without task-specific fine-tuning, completing consultations within 12.7 rounds for $T_{2}$ and 16.9 rounds for $T_{3}$.
Conclusions: The hierarchical multi-agent framework successfully enables proactive, structured medical inquiry through intelligent coordination, significantly outperforming conventional passive approaches. The model-agnostic architecture maintains high performance across different foundation models while preserving data privacy through local deployment. These findings demonstrate the potential for autonomous AI systems to enhance pre-consultation efficiency and quality in clinical settings.
@@
[…]
@@ 1 Introduction
The global healthcare system confronts unprecedented challenges from escalating patient volumes, aging populations, and persistent imbalances between medical demand and workforce capacity. This crisis manifests particularly in consultation durations: systematic reviews across 67 countries reveal that primary care visits last 5 minutes or less in nations representing half the world’s population [10]. In China, consultation times as brief as 4.3 minutes have been documented [21]—durations insufficient for ensuring diagnostic quality and comprehensive patient care. These temporal constraints not only elevate physician burnout risks but fundamentally compromise healthcare delivery quality [11, 16].
In this paper, we focus on pre-consultation, which we define as the process of guiding patients toward appropriate clinical departments (Triage) and systematically collecting essential medical history (History Taking) prior to formal consultation. Pre-consultation has emerged as a critical workflow innovation to alleviate systemic pressures: it enhances diagnostic efficiency, shortens consultation times, and reduces unnecessary clinical visits. However, traditional manual navigation approaches merely redistribute workload from physicians to nursing staff without reducing overall system burden, while also suffering from inconsistent documentation quality and fatigue-induced errors [26, 8].
Recent advances in large language models (LLMs) have enabled the development of intelligent pre-consultation systems with sophisticated dialogue capabilities. Multi-agent frameworks have demonstrated particular promise in healthcare applications, with specialized agents collaborating to enhance diagnostic reasoning [9] and achieving high accuracy in clinical advice provision [1]. However, existing systems exhibit three fundamental limitations: (1) passive interaction paradigms that respond to patient inputs rather than proactively guiding structured inquiry, failing to replicate physician-directed questioning essential for diagnostic quality; (2) context window limitations causing up to 39% performance degradation in extended conversations due to “loss-in-middle-turns” phenomena [12], compromising the 10–20 conversational exchanges required for comprehensive medical evaluation; and (3) absence of hierarchical control mechanisms preventing balanced inquiry management across multiple medical domains, resulting in systems that either over-focus on details or maintain rigid questioning sequences.
@@
To address these challenges, we propose a novel hierarchical multi-agent framework that transforms passive medical AI systems into proactive inquiry agents through autonomous task orchestration and intelligent coordination. Our approach introduces a central Controller that coordinates specialized agents to achieve three core innovations: (1) dynamic subtask completion assessment that monitors information gathering progress across 13 predefined medical domains using medically-informed evaluation criteria, (2) adaptive prompt generation that synthesizes patient responses with accumulated consultation context to formulate clinically optimal follow-up questions, and (3) hierarchical task management that balances macroscopic diagnostic progression with microscopic symptom detail collection through priority-based orchestration.
The framework formalizes the pre-consultation system into four primary tasks: Triage ($T_{1}$), History of Present Illness collection ($T_{2}$), Past History collection ($T_{3}$), and Chief Complaint generation ($T_{4}$). Among them, $T_{1}$, $T_{2}$, and $T_{3}$ are further decomposed into several domain-specific subtasks (see the table below) to support structured and comprehensive information acquisition. Through the collaboration of multiple agents—including the Monitor, Inquirer, and Controller—the system establishes an intelligent navigation framework that integrates authentic clinical logic with autonomous operational capability. The main contributions of this paper are as follows:
  • We introduce the first hierarchical multi-agent architecture with autonomous task orchestration for pre-consultation, featuring a Controller-based coordination mechanism that enables proactive, structured medical inquiry through intelligent management of specialized agent interactions.
  • We demonstrate superior performance in real-world pre-consultation scenarios, achieving high clinical quality scores across multiple evaluation dimensions on 1,372 clinical cases: Chief Complaint Similarity (4.25), History of Present Illness Similarity (3.71), and Past History Similarity (3.99), significantly outperforming conventional passive inquiry approaches while maintaining workflow coherence.
  • We establish that intelligent coordination mechanisms effectively preserve data privacy and mitigate context limitations in extended dialogues, with our locally deployed framework achieving efficient task completion ($T_{2}$: $12.7$ rounds, $T_{3}$: $16.9$ rounds) and demonstrating 98.2% information collection completeness through adaptive orchestration rather than exhaustive questioning.
Table 1: Statement of Significance
Component Description
Problem Medical consultations globally average under 5 minutes, compromising diagnostic quality. Current AI pre-consultation systems remain passive and struggle with extended dialogue management.
What is Already Known Multi-agent frameworks show promise in healthcare applications, but existing systems lack proactive inquiry capabilities and suffer from context degradation in multi-turn conversations.
What this Paper Adds We introduce the first hierarchical multi-agent architecture with autonomous task orchestration, achieving 98.2% task completion through intelligent coordination and demonstrating superior clinical quality (4.25–4.69/5.0) across 1,372 real cases.
Impact and Implications Healthcare institutions can implement this framework to reduce physician workload while maintaining consultation quality. Researchers gain a model-agnostic architecture for developing proactive medical AI systems. Patients benefit from more thorough pre-consultation assessments.
2 Related Work
2.1 Evolution of Intelligent Agents
The development of intelligent agents based on large language models (LLMs) has undergone rapid evolution from the Transformer architecture [19] replacing recurrence with attention mechanisms, enabling BERT [5] to demonstrate bidirectional context understanding and GPT-3 [3] to achieve emergent few-shot capabilities through large-scale autoregressive pretraining, with subsequent advances in GPT-4 and Gemini [15, 4] showcasing sophisticated reasoning and tool-use abilities. Recent focus has shifted toward collaborative multi-agent systems, with Liu et al. [14] and errag, Tihanyi, and Debbah [6] systematically reviewing LLM-based multi-agent frameworks across domains, while contemporary systems like MegaAgent [22] and AgentNet [25] have introduced decentralized coordination, environmental perception, and long-horizon planning, demonstrating autonomous multi-agent capabilities for complex applications.
2.2 Applications of Intelligent Agents in pre-consultation
Pre-consultation systems have evolved from rule-based dialogue trees to sophisticated multi-agent frameworks, beginning with early work by Bickmore et al.[2] establishing basic conversational patient information collection, Lewis et al.[13] improving adaptability through retrieval-augmented mechanisms integrating patient historical records, Wang and Wong [20] pioneering reinforcement learning-based multi-agent systems with specialized subtask agents, and Tang et al. [18] enhancing clinical fidelity through LLM-powered role-playing agents. Recent advances focus on coordination and contextual management with Prasad et al.[17] introducing on-demand task decomposition, Wang et al.[24] developing priority-based scheduling with dynamic agent generation, and Wang et al.[23] designing ConsultationFlow with modular specialized agents, however existing systems remain fundamentally reactive, lacking proactive inquiry capabilities and struggling with global consistency in complex medical scenarios.
3 Dataset
3.1 Data Collection and Sources
@@
We constructed a comprehensive dataset for medical triage classification by collecting Electronic Health Records (EHRs) from iiyi.com [^1] https://bingli.iiyi.com/, a well-established public medical platform with over 20 years of operational history. This platform serves as a comprehensive EHR sharing repository, encompassing electronic medical records from diverse hospitals across different regions in China. The platform maintains strict privacy standards by applying proper de-identification processes to all shared EHR data, ensuring complete protection of patient privacy information.
@@
Using systematic web crawling techniques, we initially collected 8,172 electronic health records from the platform. To ensure temporal relevance and reflect current medical practices, all collected records were sourced exclusively from the past five years. This temporal constraint guarantees that our dataset captures contemporary medical terminology, treatment approaches, and diagnostic patterns that remain clinically relevant.
3.2 Data Processing and Quality Assurance
@@
We implemented a rigorous two-stage data processing pipeline to ensure high-quality annotations and eliminate inconsistencies. The first stage involved comprehensive quality filtering where we systematically examined all essential fields including History of Present Illness (HPI), Past History (PH), and Chief Complaint (CC). Records were excluded if they contained empty fields, obvious data entry errors, abnormally short HPI descriptions, or cases where HPI and PH content demonstrated excessive similarity that could indicate data duplication or error. This stringent filtering process reduced our initial collection from 8,172 to 2,263 records.
@@
The second stage employed a hybrid validation approach combining automated prediction with expert medical review. We first applied our triage prediction model to each remaining record using the HPI, PH, and CC information as input features. Records where the predicted department assignment matched the original ground truth were automatically validated and retained. For cases where automated prediction disagreed with the original triage decision, we engaged qualified medical professionals to conduct thorough manual verification, ensuring the accuracy and reliability of our ground truth labels.
@@
Our final curated dataset comprises 1,372 validated medical records organized in a hierarchical structure reflecting real-world medical triage systems. The dataset encompasses 9 primary medical departments and 35 secondary sub-specialties, providing comprehensive coverage for both coarse-grained and fine-grained classification tasks. Figure 7 illustrates the distribution of cases across the primary medical departments.
Figure 1: Hierarchical multi-agent framework architecture for medical consultation workflow. Parts 1-3 correspond to Sec. 4.2, 4.3, 4.4 respectively.
4 Agent System
4.1 Framework Overview
@@
In this study, the pre-consultation process $T$ is conceptualized as comprising two stages: Triage and History Taking. The latter, given its large volume of information, heterogeneity of content, and the requirement for structuring doctor–patient communication, can be further organized into three modules: Chief Complaint(CC), History of Present Illness(HPI), and Past History(PH) [7].To achieve systematic information collection and optimization, we formalize the pre-consultation process $T$ as a hierarchical optimization structure comprising four main task groups: Triage ($T_{1}$), HPI collection ($T_{2}$), PH collection ($T_{3}$), and CC generation ($T_{4}$). This hierarchical modeling approach facilitates a balance between global optimality and efficient completion of individual tasks.

### v2

@@ Abstract
Methods: We developed an eight-agent architecture with centralized control mechanisms for dynamic medical consultation. The framework decomposes pre-consultation into four primary tasks: Triage ($T_{1}$), History of Present Illness collection ($T_{2}$), Past History collection ($T_{3}$), and Chief Complaint generation ($T_{4}$), with $T_{1}$$T_{3}$ further divided into 13 domain-specific subtasks. A Controller agent coordinates specialized agents through dynamic completion assessment, adaptive prompt generation, and hierarchical task management. The system was evaluated on 1,372 de-identified electronic health records a Chinese medical platform, with performance assessed across multiple foundation models (GPT-OSS 20B, Qwen3-8B, Phi4-14B). Clinical quality was independently rated by 18 licensed physicians using standardized criteria.
Results: The framework achieved 87.0% accuracy for primary department triage and 80.5% for secondary department classification. Task completion rates reached 98.2% using agent-driven scheduling versus 93.1% with sequential processing. Physician evaluation (n=18) yielded mean scores of 4.56 for Chief Complaints, 4.48 for History of Present Illness, and 4.69 for Past History on a 5-point scale, with low inter-rater variability ($\mathrm{SD}<0.23$). The system demonstrated robust cross-model generalizability without task-specific fine-tuning.
Conclusions: The hierarchical multi-agent framework demonstrates feasibility in transforming passive pre-consultation systems into proactive, physician-like inquiry agents. The model-agnostic architecture supports flexible deployment while maintaining clinical quality standards. While validation on Chinese-language records limits immediate generalizability, these findings suggest a promising pathway for enhancing pre-consultation efficiency. Further prospective validation with real patient interactions is warranted.
@@
Hierarchical multi-agent systems, Medical pre-consultation, Autonomous task orchestration, Large language models
@@ 1 Introduction
The global healthcare system faces a critical challenge: consultation times are often insufficient for comprehensive patient assessment. Systematic reviews across 67 countries reveal that primary care visits last 5 minutes or less in nations representing half the world’s population [irving2017international], with some regions documenting consultations as brief as 4.3 minutes [wang2022consultation]. These temporal constraints not only elevate physician burnout but fundamentally compromise diagnostic quality and patient safety [malene2024aging, pooyan2018resource]. Pre-consultation—defined as the systematic process of guiding patients toward appropriate departments (triage) and collecting essential medical history prior to the formal encounter—has emerged as a vital workflow innovation to alleviate these systemic pressures [zhakhina2023pre, DeGroot2022doc].
Recent advances in large language models (LLMs) have catalyzed the development of conversational AI for automated patient information collection [bickmore2015consent, tang2024medagents]. Early systems primarily utilized rule-based dialogue trees or retrieval-augmented mechanisms to improve response adaptability [lewis2021retrieval]. More recently, the field has shifted toward collaborative multi-agent architectures to handle complex clinical tasks, with frameworks demonstrating capabilities in on-demand task decomposition [Archiki2023adapt] and modular coordination among specialized agents [wang2024tdag, wang2025consultation]. However, existing approaches exhibit a fundamental limitation: they predominantly employ passive interaction paradigms that respond to patient inputs rather than proactively guiding structured inquiry. This reactive design fails to replicate physician-directed questioning—a cornerstone of diagnostic quality where clinicians systematically explore symptoms, assess severity, and rule out differential diagnoses through deliberate inquiry sequences [gawlik2024historytaking]. Furthermore, even sophisticated multi-agent architectures struggle with extended medical dialogues, experiencing up to 39% performance degradation due to context window limitations and the ”loss-in-middle” phenomenon [laban2025llms]—a critical deficiency given that comprehensive medical evaluations typically require 10–20 conversational exchanges.
We hypothesized that a hierarchical multi-agent framework with autonomous task orchestration, building upon decentralized coordination principles [xiao2025agent], could transform passive medical AI systems into proactive inquiry agents capable of physician-like structured questioning. To test this hypothesis, we pursued three primary objectives: (i) to design and implement a hierarchical multi-agent framework incorporating dynamic subtask completion assessment and adaptive prompt generation for pre-consultation; (ii) to formalize the pre-consultation workflow into structured primary tasks—Triage ($T_{1}$), History of Present Illness (HPI) collection ($T_{2}$), Past History (PH) collection ($T_{3}$), and Chief Complaint (CC) generation ($T_{4}$)—with domain-specific subtask decomposition; and (iii) to evaluate the framework’s performance in triage accuracy, task completion.
@@
To achieve these objectives, we developed a hierarchical multi-agent framework coordinating eight specialized agents through a central Controller. The framework introduces three core innovations addressing the identified limitations: (1) dynamic subtask completion assessment enabling a Monitor agent to evaluate information gathering progress across 13 predefined medical domains using clinically-informed criteria; (2) adaptive prompt generation where a Prompter agent synthesizes patient responses with accumulated context to formulate targeted, proactive follow-up questions; and (3) hierarchical task management where a Controller agent balances macroscopic diagnostic progression with microscopic symptom detail collection through priority-based orchestration. This architecture maintains contextual coherence across extended dialogues while ensuring an assertive inquiry style. We evaluated the framework using 1,372 validated medical records, assessing triage accuracy, cross-model generalizability, and physician-rated documentation quality to ensure clinical utility.
2 Methods
This study presents the development and evaluation of a hierarchical multi-agent framework for automated pre-consultation in healthcare settings. The methodology encompasses dataset construction, system architecture design, and evaluation procedures.
2.1 Study Setting and Dataset Construction
2.1.1 Data Sources
@@
We constructed a comprehensive dataset for medical triage classification by collecting Electronic Health Records (EHRs) from iiyi.com [^1] https://bingli.iiyi.com/, a well-established public medical platform with over 20 years of operational history in China. This platform serves as a large-scale online forum for sharing de-identified electronic medical records contributed by users from diverse hospitals across different regions. The platform maintains strict privacy standards by applying proper de-identification processes to all shared EHR data, ensuring complete protection of patient privacy information.
@@
Using systematic web crawling techniques, we initially collected 8,172 electronic health records from the platform. To ensure temporal relevance and reflect current medical practices, all collected records were sourced exclusively from the past five years, guaranteeing that our dataset captures contemporary medical terminology, treatment approaches, and diagnostic patterns.
2.1.2 Data Processing and Quality Assurance
We implemented a rigorous two-stage data processing pipeline to ensure high-quality annotations and eliminate inconsistencies.
@@ 2.1.2 Data Processing and Quality Assurance
The first stage involved comprehensive quality filtering where we systematically examined all essential fields including History of Present Illness (HPI), Past History (PH), and Chief Complaint (CC). Records were excluded if they met any of the following criteria: (1) empty or missing fields, (2) obvious data entry errors, (3) abnormally short HPI descriptions, or (4) excessive similarity between HPI and PH content indicating potential data duplication. This filtering process reduced our initial collection from 8,172 to 2,263 records.
@@
The second stage employed a hybrid validation approach combining automated prediction with expert medical review. We first applied our triage prediction model to each remaining record using HPI, PH, and CC information as input features. Records where the predicted department assignment matched the original ground truth were automatically validated and retained. For cases where automated prediction disagreed with the original triage decision, qualified medical professionals conducted thorough manual verification to ensure accuracy and reliability of ground truth labels.
@@
Our final curated dataset comprises 1,372 validated medical records organized in a hierarchical structure reflecting real-world medical triage systems. The dataset encompasses 9 primary medical departments and 35 secondary sub-specialties, providing comprehensive coverage for both coarse-grained and fine-grained classification tasks. Figure 1 illustrates the distribution of cases across primary medical departments.
2.2 The Hierarchical Multi-Agent Framework
2.2.1 Task Formalization
@@
We formalize the pre-consultation process $T$ as comprising two stages: Triage and History Taking. Given its large volume of information, heterogeneity of content, and the requirement for structuring doctor–patient communication, History Taking is further organized into three modules: Chief Complaint (CC), History of Present Illness (HPI), and Past History (PH) [gawlik2024historytaking].
The pre-consultation process is modeled as a hierarchical optimization structure comprising four main task groups: